OpenStudy (annipuppi):

solve the proportion x-1/32=6/24 and show work.

OpenStudy (shaik0124):

add 1/32 on both sides x-1/32+1/32=6/24+1/32 x=6/24+1/32 x=calculate u will get answer

OpenStudy (annipuppi):

so 9/32?

OpenStudy (lucaz):

\[\large{x-\frac{1}{32}~or~\frac{x-1}{32}}\]